# WASUP: Interpretable Classification with Weight-Input Alignment and Class-Discriminative SUPports Vectors

### v1

@@ Abstract
The deployment of deep learning models in critical domains necessitates a balance between high accuracy and interpretability. We introduce WASUP, an inherently interpretable neural network that provides local and global explanations of its decision-making process. We prove that these explanations are faithful by fulfilling established axioms for explanations. Leveraging the concept of case-based reasoning, WASUP extracts class-representative support vectors from training images, ensuring they capture relevant features while suppressing irrelevant ones. Classification decisions are made by calculating and aggregating similarity scores between these support vectors and the input’s latent feature vector. We employ B-Cos transformations, which align model weights with inputs to enable faithful mappings of latent features back to the input space, facilitating local explanations in addition to global explanations of case-based reasoning. We evaluate WASUP on three tasks: fine-grained classification on Stanford Dogs, multi-label classification on Pascal VOC, and pathology detection on the RSNA dataset. Results indicate that WASUP not only achieves competitive accuracy compared to state-of-the-art black-box models but also offers insightful explanations verified through theoretical analysis. Our findings underscore WASUP’s potential for applications where understanding model decisions is as critical as the decisions themselves.
Figure 1: WASUP trains a neural network to extract class-representative support feature vectors (red and blue) from training images and classifies test images by computing the similarities to its latent feature vector (white). Its B-cos architecture permits the computation of faithful local and global explanations. Stars and circles denote samples of different classes.
@@ 1 Introduction
Deep learning is increasingly becoming part of everyday life, e.g., targeted advertisements, gaming, and high-stakes areas like banking and medicine. In the latter, identifying errors in the model’s decision-making process is crucial to detect failure cases, which can be addressed by adding transparency to model prediction using explainable AI (XAI). Explanations are typically produced post-hoc for black-box models using approximation methods, which estimate the contribution of each input feature to the prediction (Bach et al., 2015; Ribeiro et al., 2016; Selvaraju et al., 2017; Lundberg & Lee, 2017; Sundararajan et al., 2017; Petsiuk, 2018). While these methods can accurately compute feature contributions for low-dimensional inputs Lundberg & Lee (2017), they introduce approximation errors for high-dimensional inputs such as images (Adebayo et al., 2018).
@@
As an alternative, inherently interpretable models provide explanations by design. For instance, logistic regression can be considered an inherently interpretable model since evaluating its coefficients directly explains the effect of each input feature towards the prediction. Translating these properties to complex functions implemented by neural networks was further fueled by the argument that if a technique is able to summarize the decision-making of a black box without any error, the black box could be replaced by an inherently interpretable model in the first place (Rudin, 2019). As a result, a growing research community is dedicated to developing inherently interpretable models, e.g., for decision-critical tasks like medicine Barnett et al. (2021); Kim et al. (2021); Wolf et al. (2023).
@@
The explanations provided by these models can be categorized into local and global explanations, and evaluated in terms of axioms to verify if they are faithful Sundararajan et al. (2017)  [^1] Throughout this manuscript, we define faithful explanations as explanations that satisfy all of the axioms proposed by Sundararajan et al. (2017): Completeness, Sensitivity, Implementation Invariance, Dummy, Linearity, and Symmetry-Preserving.. Local explanations focus on individual predictions, revealing how the model arrived at a specific outcome for a particular instance. In contrast, global explanations provide insight into the model’s overall behavior. In logistic regression, global explanations are derived from the model’s coefficients, which describe how each feature influences the prediction across all instances. To produce local explanations, a sample’s specific feature values are multiplied by their corresponding coefficients, illustrating how each feature contributed to the particular prediction.
@@
As outlined in Tab. 1, locally interpretable deep learning models include the work by Wang & Sabuncu (2022), who trained a non-parametric classification head with metric learning to classify an image by summing the softmax-normalized similarity scores between training image latent feature vectors and the test image latent feature vector. They proposed to provide local explanations by showing the most influential images alongside their corresponding scores for a given test sample, similar to the k-nearest-neighbor algorithm. While intuitive, the approach lacks contribution maps in pixel space to identify important input features. To this end, BagNets (Brendel & Bethge, 2019) classify local patches of an image individually and sum the probabilities of all patches for the final classification. Providing the class-probability maps of the individual patches’ predictions as explanations shows which local structures showed evidence of a respective class. However, they lack faithful fine-grained attributions computed on a pixel-level, which are provided by B-cos networks (Böhle et al., 2022; 2024). In these models, each forward pass is summarized as a linear transformation, and its resulting weight matrix then serves to compute local explanations in image space. More recently, they were used to guide knowledge distillation by Parchami-Araghi et al. (2024), who showed that a student network trained with knowledge distillation detects the same features as the teacher network, which explanation-guiding helped to overcome. However, B-cos models lack global interpretability because the weight matrix is dynamically computed for each input.
Table 1: Comparison of inherently interpretable models and their properties.
Model Local Expl. Global Expl. Faithful Expl.
NW-Head Wang & Sabuncu (2022)
ProtoPNet Chen et al. (2019)
BagNets Brendel & Bethge (2019)
B-Cos Böhle et al. (2022)
WASUP (ours)
@@
Models that provide global interpretability include concept bottleneck models (Koh et al., 2020), which train an encoder to predict concepts present in an image and a classifier to predict the target class from these concepts, which serve as local explanations during a prediction. However, they require access to concepts, i.e., meta information on each concept present in an image, e.g., the color of the tail, legs, or head of a bird. While the concepts themselves provide global and local explanations and are manually pre-defined, concept bottleneck models lack explanations in the image domain like the Nadaraya-Watson head (Wang & Sabuncu, 2022). The most prominent inherently interpretable model is ProtoPNet (Chen et al., 2019) and its extensions (Kim et al., 2021; Donnelly et al., 2022), which provide local and global explanations. Its key idea is to learn class-specific prototypical parts across the training set to facilitate reasoning as “this part of an unseen test image looks like that part from my training images.” ProtoPNet implements its reasoning by computing similarity scores between learned latent feature vectors and (test sample) latent feature vectors. To provide image space explanations, it upsamples similarity maps from latent space into the input domain. However, Hoffmann et al. (2021) found that the explanations provided by ProtoPNet can be misleading and called for more rigorous evaluations.  Sacha et al. (2024) proposed a benchmark to evaluate spatial misalignment of explanations given by ProtoPNet, which was also found by the theoretical evaluation of Wolf et al. (2024). They showed that spatial relationship between the input space and latent space can not be guaranteed for ProtoPNets with deep convolutional neural networks encoders, generalizing to other prototypical networks like Pip-Net (Nauta et al., 2023), ProtoTrees (Nauta et al., 2021), XProtoNet (Kim et al., 2021), or the Deformable ProtoPNet (Donnelly et al., 2022). While prototypical part-based case-based reasoning has shown to be the most human comprehensible explanation (Nguyen et al., 2021; Kim et al., 2023; Jeyakumar et al., 2020), the number of explanations to be evaluated in its current implementations grows with the number of prototypical parts per class learned by the model.
@@
We propose WASUP, an inherently interpretable neural network that fulfills all three properties: local, global, and faithful explanations (see Tab. 1). We inject case-based reasoning into WASUP’s decision-making to provide local and global explanations that reduce the complexity of explanations to be evaluated. WASUP can implement any neural network architecture (e.g., ResNets (He et al., 2016), Transformers (Vaswani, 2017), Mamba (Gu & Dao, 2023)) in its feature extractor without limiting their computational capacity. Specifically, the model extracts support vectors from training images that capture only class-representative features by suppressing negative contributions from other classes. Classification is performed by computing and summing similarity scores between these support vectors and the latent feature vector of the input image (see Fig. 1). Levering B-cos networks ensures that both the support vectors and their similarity to test images are directly and faithfully explainable in the input space, enabling an intuitive understanding of the model’s reasoning. In summary, the support samples serve two purposes: (i) they help the model capture the high variability between different classes in the data, and (ii) they provide the user with examples that illustrate which parts of the class-representative training images the feature extractor focuses on. Our contributions are as follows:
@@
  • We propose WASUP, an inherently interpretable network that provides faithful local and global explanations for image classification with either single or multiple labels.
@@
  • We prove that explanations provided by WASUP fulfill the axioms required to be faithful.
@@
  • We empirically evaluate WASUP and its explanations on three tasks (fine-grained single-label image classification on Stanford Dogs, pathology prediction on RSNA, multi-label image classification on Pascal VOC), and three architectures with different numbers of parameters (DenseNet121: 8M, ResNet50: 26M, a hybrid vision transformer: 81M).
  • We demonstrate how the explanations provided by WASUP enable model debugging.
@@ 2 Background
Wang & Sabuncu (2022) proposed to classify images with the non-parametric Nadaraya-Watson head and a convolutional neural network $\mathcal{F}_{\theta}$ with parameters $\theta$ extracting latent vector of dimension $d$. During training, they randomly sample a number of support vectors $v_{k}^{c}$ for every batch per class from the training set $\{(\mathcal{I}_{k},y_{k})\}_{k=1}^{N}$, $V^{c}=\left\{v_{k}^{c}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\right\}$, with $\Omega(V^{c})=N_{s}$. This selection is executed for every training batch, thereby directing the optimization of the feature extractor $\mathcal{F}_{\theta}$ to form distinct, class-specific clusters within the latent feature space.
@@
Local explanations show the images of the most influential support samples in terms of softmax probabilities (see Fig. 2). At first glance, these explanations seem intuitive, but the main limitations are two-fold: (i) Computing probabilities from (inverted) Euclidean distances with softmax: Regardless of the input (e.g., in an adversarial attack), the model explanations will suggest that at least one support sample is similar with sufficient probability. Additionally, the softmax probabilities are inapplicable to multi-label classification tasks out of the box since they sum to one. Thus, the model cannot predict three classes simultaneously under a standard decision threshold of 0.5. (ii) No pixel-level explanations: While explanations using similar input images are intuitive for humans, the lack of attribution maps hinders a further understanding of the network’s mechanisms and the detection of biased predictions.
@@
B-cos networks proposed by Böhle et al. (2022) impose weight-input alignment between the layer weights and input by re-formulating the scalar product used by neural network computations, removing biases and standard non-linearities. This forces the network to focus on the most salient features across the training set and facilitates the summary of each forward pass as an input-dependent linear equation. Each input feature’s contribution can directly be traced by its corresponding weight in the corresponding transformation matrix.
Mathematically, the B-cos transformation computes a modified dot-product, that increases weight-input alignment its exponent $B$, between an input vector $x$ and a layer weight $w$:
@@
with $\hat{w}=w/\lVert w\rVert\implies\lVert\hat{w}\rVert=1$, cos the cosine similarity, sgn the sign function, and $\times$ the real valued multiplication. While it is bounded by the magnitude of $\|x\|$, it effectively computes an input-dependent linear equation, in matrix form expressed by:
@@
where $\lvert cos(x,\mathbf{\hat{W}})\rvert^{B-1}$ scales the elements of the fixed weight matrix $\hat{\mathbf{W}}$ ($\odot$ the elementwise multiplication). A neural network $\mathcal{F}_{\theta}$ with parameters $\theta$ that computes a series of these equations is called piece-wise linear and its forward pass is summarized by the single input-dependent weight matrix:
$\mathcal{F}_{\theta}({x})=\tilde{\mathbf{W}}_{L}(a_{L})\tilde{\mathbf{W}}_{L-1}(a_{L-1})\dots\tilde{\mathbf{W}}_{1}(a_{1}=x)x=\left(\prod_{j=1}^{L}\tilde{\mathbf{W}_{j}}(a_{j})\right)x=\mathbf{W}_{1\rightarrow L}(x)x.$ (3)
Thus, the contribution maps in terms of a pixel location $(m,n)$ in image space for a single forward pass are faithfully computed with:
$\phi^{l}_{j}(x)_{(m,n)}=\sum_{ch}\left[\left[\mathbf{W}_{1\rightarrow l}\right]^{T}_{j}\odot x\right]_{(ch,m,n)},$ (4)
with $l$ any layer and $j$ the index of the row (neuron) in $\mathbf{W}_{1\rightarrow l}$ (note that this allows to explain any intermediate layer’s $l$ neurons $j$). Since B-cos networks are trained with binary cross-entropy loss, features negatively attributing the class logit cannot increase its probability and are not visualized.
Böhle et al. (2022) propose to compute RGBA explanations by normalizing each color channel [ch, 1-ch] [^2] B-cos networks require an image encoding of [R, G, B, 1-R, 1-G, 1-B] in the channel dimension to uniquely encode colors and to mitigate favoring of brighter regions. to sum to 1 to maintain the angle of each color pair and to then scale RGBA values into the range [0..1]. Finally, the alpha value (pixel opacity) is computed with the 99.9th percentile $p_{99.9}$ of the resulting weight $w$ as $\text{min}\left(\left[\lVert w_{(m,n)}\rVert_{2}\right]/\left[p_{99.9}(\lVert w_{(m,n)}\rVert_{2})\right],1\right)$.
The obvious strength of B-cos models is that they are able to extract local explanations of a single forward pass that resemble the transformation of an input image with the network with high fidelity, which we prove to fulfill the axioms required for faithful explanations in the appendix A.1. Nonetheless, they do not provide global explanations (see Fig. 2).
Figure 2: While the Nadaraya-Watson head Wang & Sabuncu (2022) provides explanations following case-based reasoning, and B-cos networks Böhle et al. (2022) faithfully explain the image transformation of a neural network, WASUP combines both properties, thereby adding global explanations by extracting a fixed set of class-representative, explainable support vectors.
@@ 3 Methods
We propose WASUP, which combines the strengths of B-cos networks and the Nadaraya-Watson head, to obtain an inherently interpretable neural network that implements case-based reasoning and provides faithful local and global explanations, thereby reducing the complexity of explanations to evaluate. WASUP achieves this by extracting class-representative support vectors from training images for similarity-based classification and by leveraging the B-cos transform to faithfully explain its decision, see Fig. 2.

### v2

# SIC: Similarity-Based Interpretable Image Classification with Neural Networks

@@ Abstract
The deployment of deep learning models in critical domains necessitates a balance between high accuracy and interpretability. We introduce SIC, an inherently interpretable neural network that provides local and global explanations of its decision-making process. Leveraging the concept of case-based reasoning, SIC extracts class-representative support vectors from training images, ensuring they capture relevant features while suppressing irrelevant ones. Classification decisions are made by calculating and aggregating similarity scores between these support vectors and the input’s latent feature vector. We employ B-Cos transformations, which align model weights with inputs, to yield coherent pixel-level explanations in addition to global explanations of case-based reasoning. We evaluate SIC on three tasks: fine-grained classification on Stanford Dogs and FunnyBirds, multi-label classification on Pascal VOC, and pathology detection on the RSNA dataset. Results indicate that SIC not only achieves competitive accuracy compared to state-of-the-art black-box and inherently interpretable models but also offers insightful explanations verified through practical evaluation on the FunnyBirds benchmark. Our theoretical analysis proves that these explanations fulfill established axioms for explanations. Our findings underscore SIC’s potential for applications where understanding model decisions is as critical as the decisions themselves.
@@ 1 Introduction
Deep learning is increasingly becoming part of everyday life, e.g., targeted advertisements, gaming, and high-stakes areas like banking and medicine. In the latter, identifying errors in the model’s decision-making process is crucial to detect failure cases, which can be addressed by adding transparency to model predictions using explainable AI (XAI). Explanations are typically produced post-hoc for black-box models using approximation methods, which estimate the contribution of each input feature to the prediction [3, 24, 28, 19, 30, 23]. While these methods can accurately compute feature contributions for low-dimensional inputs [19], they introduce approximation errors for high-dimensional inputs such as images [1].
@@
As an alternative, inherently interpretable models provide explanations by design. For instance, logistic regression can be considered an inherently interpretable model since evaluating its coefficients directly explains the effect of each input feature towards the prediction. Translating these properties to complex functions implemented by neural networks was further fueled by the argument that if a technique is able to summarize the decision-making of a black box without any error, the black box could be replaced by an inherently interpretable model in the first place [26]. As a result, a growing research community is dedicated to developing inherently interpretable models, e.g., for decision-critical tasks like medicine [4, 16, 32].
@@
The explanations provided by these models can be categorized into local and global explanations, and evaluated in terms of axioms to verify if they are faithful [30]  [^1] Throughout this manuscript, we define faithful explanations as explanations that satisfy all of the axioms proposed by Sundararajan et al. [30]: Completeness, Sensitivity, Implementation Invariance, Dummy, Linearity, and Symmetry-Preserving.. Local explanations focus on individual predictions, revealing how the model arrived at a specific outcome for a particular instance. In contrast, global explanations provide insight into the model’s overall behavior. In logistic regression, global explanations are derived from the model’s coefficients, which describe how each feature influences the prediction across all instances. To produce local explanations, a sample’s specific feature values are multiplied by their corresponding coefficients, illustrating how each feature contributed to the particular prediction.
Related Work
@@
As outlined in Tab. 1, locally interpretable deep learning models include the work by Wang and Sabuncu [31], who trained a non-parametric classification head with metric learning to classify an image by summing the softmax-normalized similarity scores between training image latent feature vectors and the test image latent feature vector. They proposed to provide local explanations by showing the most influential images alongside their corresponding scores for a given test sample, similar to the k-nearest-neighbor algorithm. While intuitive, the approach lacks contribution maps in pixel space to identify important input features. To this end, BagNet [6] classify local patches of an image individually and sum the probabilities of all patches for the final classification. Providing the class-probability maps of the individual patches’ predictions as explanations shows which local structures showed evidence of a respective class. However, they lack faithful fine-grained attributions computed on a pixel-level, which are provided by B-cos networks [5, 7]. In these models, each forward pass is summarized as a linear transformation, and its resulting weight matrix then serves to compute local explanations in image space. However, B-cos models lack global interpretability because the weight matrix is dynamically computed for each input.
Model Local Global Faithful
NW-Head [31]
ProtoPNet [8]
BagNet [6]
B-Cos [5] (✓)
SIC (ours)
Table 1: Comparison of the ability of inherently interpretable models to provide local, global, and faithful explanations.
@@
Models that provide global interpretability include concept bottleneck models [18], which train an encoder to predict concepts present in an image and a classifier to predict the target class from these concepts, which serve as local explanations during a prediction. However, they require access to concepts, i.e., meta information on each concept present in an image, e.g., the color of the tail, legs, or head of a bird. While the concepts themselves provide global and local explanations and are manually pre-defined, concept bottleneck models lack explanations in the image domain. The most prominent inherently interpretable model is ProtoPNet [8] and its extensions [16, 9], which provide local and global explanations. Its key idea is to learn class-specific prototypical parts across the training set to facilitate reasoning as “this part of an unseen test image looks like that part from my training images.” To provide image space explanations, it upsamples distance maps from the latent space into the input domain. However, Hoffmann et al. [13] found that the explanations provided by ProtoPNet can be misleading and called for more rigorous evaluations.  Sacha et al. [27] proposed a benchmark to evaluate spatial misalignment of explanations given by ProtoPNet, which was also found by the theoretical evaluation of Wolf et al. [33]. They showed that spatial relationship between the input space and latent space can not be guaranteed for ProtoPNets with deep convolutional neural networks encoders, generalizing to other prototypical networks like Pip-Net [21], ProtoTrees [20], XProtoNet [16], or the Deformable ProtoPNet [9]. While prototypical part-based case-based reasoning has shown to be the most human comprehensible explanation [22, 17, 15], the number of explanations to be evaluated in its current implementations grows with the number of prototypical parts per class learned by the model.
Contribution
@@
We propose SIC, a transparent neural network for similarity-based interpretable image classifciation, that fulfills all three properties: local, global, and faithful explanations (see Tab. 1). We inject case-based reasoning into SIC’s decision-making to provide local and global explanations that reduce the complexity of explanations to be evaluated. SIC can implement any neural network architecture in its feature extractor without limiting their computational capacity. Specifically, the model extracts support vectors from training images that capture only class-representative features by suppressing negative contributions from other classes. Classification is performed by computing and summing similarity scores between these support vectors and the latent feature vector of the input image (see Fig. LABEL:fig:fig1). Leveraging B-cos networks ensures that both the support vectors and their similarity to test images are directly and faithfully explainable in the input space, enabling an intuitive understanding of the model’s reasoning. In summary, the support samples serve two purposes: (i) they help the model capture the high variability between different classes in the data, and (ii) they provide the user with examples that illustrate which parts of the class-representative training images the feature extractor focuses on. Our contributions are as follows:
@@
  • We propose SIC, an inherently interpretable network that provides faithful local and global explanations for image classification with either single or multiple labels.
@@
  • We prove that explanations provided by SIC fulfill the axioms required to be faithful.
  • We practically evaluate SIC’s explanations based on the criteria proposed in the FunnyBirds framework and its associated dataset.
@@
  • We empirically evaluate SIC and its explanations on three tasks (fine-grained single-label image classification on Stanford Dogs, pathology prediction on RSNA, multi-label image classification on Pascal VOC), and three architectures with different numbers of parameters (DenseNet121: 8M, ResNet50: 26M, a hybrid vision transformer: 81M).
  • We demonstrate how the explanations provided by SIC enable model debugging.
@@ 2 Background
Wang and Sabuncu [31] proposed to classify images with the non-parametric Nadaraya-Watson head and a convolutional neural network $\mathcal{F}_{\theta}$ with parameters $\theta$ extracting latent vector of dimension $d$. During training, they randomly sample a number of support vectors $v_{k}^{c}$ for every batch per class from the training set $\{(\mathcal{I}_{k},y_{k})\}_{k=1}^{N}$, $V^{c}=\left\{v_{k}^{c}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\right\}$, with $\Omega(V^{c})=N_{s}$. This selection is executed for every training batch, thereby directing the optimization of the feature extractor $\mathcal{F}_{\theta}$ to form distinct, class-specific clusters within the latent feature space.
@@
Local explanations show the images of the most influential support samples in terms of softmax probabilities. At first glance, these explanations seem intuitive, but the main limitations are two-fold: (i) Computing probabilities from (inverted) Euclidean distances with softmax: Regardless of the input (e.g., in an adversarial attack), the model explanations will suggest that at least one support sample is similar with sufficient probability. Additionally, the softmax probabilities are inapplicable to multi-label classification tasks out of the box since they sum to one. Thus, the model cannot predict three classes simultaneously under a standard decision threshold of 0.5. (ii) No pixel-level explanations: While explanations using similar input images are intuitive for humans, the lack of attribution maps hinders a further understanding of the network’s mechanisms and the detection of biased predictions.
@@
B-cos networks proposed by Böhle et al. [5] impose weight-input alignment between the layer weights and input by re-formulating the scalar product used by neural network computations, removing biases and standard non-linearities. This forces the network to focus on the most salient features across the training set and facilitates the summary of each forward pass as an input-dependent linear equation. Each input feature’s contribution can directly be traced by its corresponding weight in the corresponding transformation matrix.
Mathematically, the B-cos transformation computes a modified dot-product, that increases weight-input alignment with an exponent $B$, between an input vector $x$ and a layer weight $w$:
@@
$\mathcal{F}_{\theta}({x})$ $=\tilde{\mathbf{W}}_{L}(a_{L})\tilde{\mathbf{W}}_{L-1}(a_{L-1})\dots\tilde{\mathbf{W}}_{1}(a_{1}=x)x$
$=\left(\prod_{j=1}^{L}\tilde{\mathbf{W}_{j}}(a_{j})\right)x=\mathbf{W}_{1\rightarrow L}(x)x.$ (1)
Thus, the contribution maps in terms of a pixel location $(m,n)$ in image space for a single forward pass are computed with $\phi^{l}_{j}(x)_{(m,n)}=\sum_{ch}\left[\left[\mathbf{W}_{1\rightarrow l}\right]^{T}_{j}\odot x\right]_{(ch,m,n)}$, with $l$ any layer and $j$ the index of the row (neuron) in $\mathbf{W}_{1\rightarrow l}$ (note that this allows to explain any intermediate layer’s $l$ neurons $j$). Since B-cos networks are trained with binary cross-entropy loss, features negatively attributing the class logit cannot increase its probability and are not visualized.
Böhle et al. [5] proposed to compute RGBA explanations by normalizing each color channel pair [^2] B-cos networks require an image encoding of [R, G, B, 1-R, 1-G, 1-B] in the channel dimension to uniquely encode colors and to mitigate favoring of brighter regions. to sum to 1 to maintain the angle of each color pair and to then scale RGBA values into the range [0..1]. The alpha value (pixel opacity) is computed with the 99.9th percentile $p_{99.9}$ of the resulting weight $w$ as $\text{min}\left(\left[\lVert w_{(m,n)}\rVert_{2}\right]/\left[p_{99.9}(\lVert w_{(m,n)}\rVert_{2})\right],1\right)$ and smoothed with an average kernel of size nine.
B-cos models offer the notable advantage of generating local explanations from a single forward pass that resembles the network’s transformation of the input image. However, whether B-cos meets the axioms of faithful explanations remains unclear. In Appendix A.1, we prove that B-cos satisfies these necessary axioms.
@@ 3 Methods
We propose SIC, which combines the strengths of B-cos networks and the Nadaraya-Watson head, to obtain an inherently interpretable neural network for similarity-based classification that provides faithful local and global explanations with a focus on minimizing the number of explanations requiring review. SIC achieves this by extracting class-representative support vectors from training images for similarity-based classification and by leveraging the B-cos transform to faithfully explain its decision.